# ShortcutBreaker: Low-Rank Noisy Bottleneck with Global Perturbation Attention for Multi-Class Unsupervised Anomaly Detection

### v1

@@ 3.3 Global Perturbation Attention (GPA)
where $Q\in\mathbb{R}^{N\times d}$, $K\in\mathbb{R}^{N\times d}$ and $V\in\mathbb{R}^{N\times d}$ indicate the query, key and value vectors respectively. $K^{T}\in\mathbb{R}^{d\times N}$ is the transpose matrix of $K$ and $d_{k}$ is the scaling factor.
However, as shown in Fig. 4 ((a) vs. (c) and (d) vs. (f)), we observe that the Softmax function tends to overfocus on dominant activation regions during output generation. This behavior is prone to inducing identity mapping under the MUAD setting (Guo et al., 2025). To avoid the overfocusing behavior, the most intuitive approach is to remove Softmax or replace it with an alternative, the Sigmoid (Eq. (6)).
$\operatorname{Attention}(Q,K,V)=\operatorname{sigmoid}\left(\frac{QK^{\top}}{\sqrt{d_{k}}}\right)V$ (6)
As shown in the 1st and 2nd columns of Fig. 4, when Softmax is removed, the overconcentration issue is alleviated. When Softmax is replaced with the Sigmoid function, the attention spreads more broadly. This indicates that Sigmoid functions tend to leverage longer-range dependencies for feature reconstruction rather than relying on focal attention, thereby reducing the likelihood of propagating unseen information to subsequent layers. Therefore, we apply Sigmoid as a replacement for Softmax in the final setting.
Global-Self Masking (GSM)
Following global redistribution, GPA applies a global-self-masking operation to the attention map to suppress shortcut learning. The self-masking mechanism effectively prevents tokens from attending to their own positions.
Figure 5: The flowchart of Global-Self Masking operation. AM: attention map and ($\cdot$) indicate element-wise multiplication.
Moreover, considering that feature tokens are captured by the ViT structure, all tokens in the same sequence may share similar information. Therefore, we propose to mask tokens globally and randomly. Specifically, we implement self-masking by masking the elements on the diagonal of the attention map (AM) and conduct random global masking via attention dropout. The flowchart of GSM is shown in Fig. 5; the output AM is obtained by element-wise multiplication of the input AM and our mask.
@@
Image-Level Pixel-Level Image-Level Pixel-Leve
@@ 4.1 Experimental Setting
Datasets
We evaluated our method on four well-established datasets, including three for industrial scenarios: MVTec-AD (Bergmann et al., 2019), ViSA (Zou et al., 2022), Real-IAD (Wang et al., 2024), and one for medical scenario: Universal Medical (He et al., 2024b; a; Zhang et al., 2023; 2024a). MVTec-AD comprises 15 object categories, with 3,629 normal images in the training set and 498 normal images alongside 1,982 anomalous samples in the test set. ViSA contains 12 categories, providing 8,659 normal training images and a test set with 962 normal images and 1,200 anomalous cases. Real-IAD, the largest industrial benchmark, includes 30 diverse objects, utilizing 36,645 normal images for training and 63,256 normal images combined with 51,329 anomalous instances for testing. Universal Medical consists of 13,339 normal cases for training, with a test set containing 2,514 normal cases and 4,499 abnormal cases. This dataset spans three medical imaging modalities: Brain MRI, Liver CT, and Retinal CT scans. All datasets include both image-level and pixel-level labels.
Metrics
For evaluation metrics, we followed protocols from (Zhang et al., 2023; He et al., 2024a; Guo et al., 2025), adopting four metrics: Area Under the Receiver Operating Curve (AUC), F1-max score, average precision (AP) for both image-level detection and pixel-level localization tasks, and Area Under the Per-Region-Overlap (AUPRO) to further evaluate the localization task.
Implementation details
In our experiments, the encoder is initialized with DINO pre-trained weights (Darcet et al., 2023) and kept frozen during training. Images are resized to 512×512 and then center-cropped to 448×448. We use a stable variant of the AdamW optimizer (Wortsman et al., 2023) incorporating the AMSGrad algorithm (Reddi et al., 2019), with a batch size of 32. Training iterations are set to 20,000 for MVTec-AD and ViSA, 25,000 for Universal Medical, and 50,000 for the largest dataset, Real-IAD. The learning rate is initialized to 2e-3 and gradually reduced to 2e-4 via a Cosine Annealing schedule with a warm-start scheme (Loshchilov & Hutter, 2016). The model is optimized using the global hard-mining loss (Guo et al., 2025). In LRNB, we set the noise rate to 0.1, and set the number of LRNB layers to 2 for MVTec-AD and ViSA, and 3 for Universal Medical and Real-IAD.
@@ 4.2 Evaluations
We evaluate our method against seven state-of-the-art (SOTA) MUAD methods—UniAD (You et al., 2022), DiAD (He et al., 2024b), ViTAD (Zhang et al., 2023), InvAD (Zhang et al., 2024b), Recontrast (Guo et al., 2024), and Dinomaly (Guo et al., 2025)—across four benchmark datasets: MVTec-AD (Bergmann et al., 2019), ViSA (Zou et al., 2022), Universal Medical (He et al., 2024b), and Real-IAD (Wang et al., 2024). Performance is quantified using both image-level metrics (I-AUC, I-AP, I-F1) and pixel-level metrics (P-AUC, P-AP, P-F1, P-AUPRO), where higher values indicate better detection capability. Experimental results are presented in Table 1, where our method outperforms comparative methods across all datasets in most metrics (The qualitative results obtained by our method can be seen in Appendix 3). On the widely adopted MVTec-AD, our method achieves SOTA overall performance, with the highest image-level metrics of 99.8/99.9/99.5, as well as three top-ranked and one second-ranked pixel-level metrics of 98.4/69.4/68.9/95.2. On ViSA, our method consistently achieves the best image-level performance of 98.9/99.1/96.2 and competitive pixel-level performance of 99.0/53.1/56.2/94.5. These results demonstrate that image-level performance on these two datasets is nearly 100On Universal Medical, our method attains the highest image-level metrics of 87.8/87.8/82.5, surpassing prior SOTAs by a large margin of 2.6/3.4/0.5, and achieves best or third-best pixel-level metrics of 97.1/54.8/54.0/87.1. On Real-IAD, our method produces a new SOTA result, with image-level and pixel-level performance of 90.6/87.9/81.4 and 99.1/44.7/49.1/95.6, outperforming previous SOTAs by 1.3/1.1/1.1 and 0.3/1.9/2.0/1.7. These results demonstrate strong generalization to complex medical and diverse real-world industrial scenarios.
@@
Table 2: Ablations studies of component contributions on Universal Medical dataset, including LRNB: low-rank noisy bottleneck, GRD: global redistribution operation and GSM: global-self masking mechanism. GPA: global perturbation attention, which is the combination of GRD and GSM. (%)
LRNB GPA Image-level Pixel-level
GRD GSM AUC AP F1 AUC AP F1 AUPRO
79.28 78.67 80.41 95.49 39.30 41.11 81.35
86.38 85.72 81.67 96.83 51.83 52.29 86.18
79.89 78.60 80.26 95.70 41.03 42.11 81.77
79.48 78.77 80.50 95.53 39.54 41.40 81.78
87.53 86.96 82.42 97.05 54.74 53.89 87.11
86.70 87.12 81.53 96.98 53.43 52.88 86.09
87.81 87.75 82.45 97.06 54.81 54.03 87.06
@@ 4.3 Ablation Studies
To explore the contributions of each component in our method, including the low-rank noisy bottleneck (LRNB), global redistribution (GRD), and global-self masking (GSM) operations, we conduct ablation experiments as shown in Table 2. The baseline is constructed following Dinomaly (Guo et al., 2025) and ViTAD (Zhang et al., 2023), which construct a baseline model with a DINO-pretrained ViT encoder and a learnable softmax attention-based ViT decoder. The effectiveness of the baseline has been proven in industrial scenarios. The results of the ablation experiments demonstrate that the proposed LRNB, GRD, and GSM modules all contribute to performance improvement, with their combinations further enhancing performance. Compared to the baseline model, using any single module alone achieves better performance in terms of the key indicators I-AUC and P-AUC. Among them, GRD and GSM bring moderate improvements, while LRNB shows the most significant enhancement—it notably boosts image-level and pixel-level performances with respective large margins of 7.10/7.05/1.26 and 1.44/12.53/11.18/4.83. These results demonstrate that the LRNB module serves as a core foundation in our method. Moreover, module combinations yield even better results than single modules. When LRNB is combined with either GRD or GSM, performance exceeds that of LRNB used alone. The combination of LRNB and GRD stands out particularly in optimizing I-AUC and P-AUC. The integration of all three modules (LRNB + GRD + GSM) delivers the best overall performance, achieving the highest image-level metrics 87.81/87.80/82.45 and top/top-2 ranked pixel-level performance 97.06/54.81/54.03/87.06—confirming that their collaborative effect effectively enhances the model’s overall performance.
BN Image-level Pixel level
AUC AP F1 AUC AP F1 AUPRO
None 78.6 77.7 80.8 95.3 38.6 40.1 81.2
FJ 77.7 74.0 80.2 95.6 39.5 42.3 81.1
NDB 82.8 81.8 80.4 96.3 44.6 46.0 83.9
LRNB 87.8 87.8 82.5 97.1 54.8 54.0 87.1
Table 3: Performance comparison of different bottlenecks (BNs) on Universal Medical. (%)
Modules Image-level Pixel level
AUC AP F1 AUC AP F1 AUPRO
ViT 86.4 85.7 81.7 96.8 51.8 52.3 86.2
CNN 84.6 84.9 81.3 96.4 48.4 48.9 85.1
CNN+ViT 83.9 83.8 81.2 96.4 47.8 48.4 84.1
NMA 87.4 86.6 82.0 97.0 53.2 53.6 86.4
GPA 87.8 87.8 82.5 97.1 54.8 54.0 87.1
Table 4: Performance comparison of different modules in decoder on Universal Medical. (%)
Comparison between LRNB and previous noisy bottlenecks
Previous methods also proposed some noisy operations in the bottleneck to avoid identity shortcuts, such as feature jittering (FJ) in UniAD (You et al., 2022) and dropout noisy bottleneck (DNB) in Dinomaly (Guo et al., 2025). To further validate the advantage of the proposed LRNB, we conduct a comparison in Table 4. The results in this table demonstrate that FJ experiences a performance drop and DNB yields slight improvements compared to models without such noise operations, while our LRNB achieves a significant enhancement.
@@
Figure 6: The plots of loss (left part) and the averaged anomaly scores (right part) of different bottlenecks on Universal Medical.
@@
To explore the capacity of these operations in addressing the ”identity mapping” issue, we further examine the corresponding training loss and average anomaly score. As shown in Fig. 6, models without a bottleneck achieve near-zero training loss and obtain extremely close anomaly scores between normal and abnormal cases (0.0195 vs. 0.0223). These results indicate that these models suffers from identity mapping behavior where inputs are directly replicated in outputs. Introducing FJ or DNB noise operations partially alleviates this issue, which we attribute to enhanced noise robustness from exposure to diverse medical patterns. Our LRNB framework effectively resolves identity mapping: while moderately increasing the normal score, it substantially enhances the abnormal score. This strategic trade-off widens the discrimination margin between normal and abnormal samples, ultimately improving anomaly detection performance.
Comparison between GPA and previous modules in decoder
Prior works proposed decoder modules to address the identity mapping issue. (Lu et al., 2024) attributes it to encoder-decoder homogeneity and advocates heterogeneous decoders for reconstructing encoder outputs (e.g., using ViT as encoder and CNN as decoder blocks). (You et al., 2022) introduced neighbor-masking attention (NMA) to prevent information leakage from CNN-extracted tokens. We compared ViT (baseline), CNN, CNN+ViT, NMA, and our GPA in Table 4. The table shows that replacing ViT with CNN or CNN+ViT yields poorer performance, as CNNs are more prone to identity mapping (You et al., 2022). Fig. 7 in Appendix 2.1 supports this: training losses of CNN and CNN+ViT drop sharply—even lower than ViT’s—narrowing the score distance between normal and abnormal cases. In contrast, NMA and our GPA outperform ViT, where the corresponding increased training losses and reconstruction errors for both cases confirm they alleviate shortcut learning. Finally, our GPA outperforms NMA in most metrics, which we attribute to its adaptability to ViT-extracted tokens.

### v2

@@ 2.3 Frequency-Filtering Block
where $Q\in\mathbb{R}^{N\times d}$, $K\in\mathbb{R}^{N\times d}$ and $V\in\mathbb{R}^{N\times d}$ are projected from input tokens $X\in\mathbb{R}^{N\times d}$, indicating the query, key and value vectors respectively. $K^{T}\in\mathbb{R}^{d\times N}$ is the transpose matrix of $K$ and $d_{k}$ is the scaling factor.
In anomaly detection, Transformer blocks have recently been adopted to replace conventional convolution blocks due to their global modeling capability (You et al., 2022; Zhang et al., 2023; Guo et al., 2025). However, the self-attention mechanism often learns a ”spiky” weight distribution, primarily focusing on the query’s own location, which facilitates an identity shortcut (You et al., 2022; Guo et al., 2025).
Motivated by three intuitive benefits, we construct a Frequency-Filtering Block (FFB) by replacing the self-attention mechanism by a Frequency-Filtering Unit, as shown in Fig. 3(c). First, the FFU processes token sequences in the frequency domain using the 2D DFT. According to the properties of the DFT, every spectral coefficient encapsulates information from all spatial locations. By performing mixing in this spectral space, the FFU inherently maintains a holistic global modeling capability, forcing the model to reconcile inputs with a global distribution. This effectively prevents the network from relying on isolated localized shortcuts for reconstruction. Second, unlike standard self-attention $A(X)=\text{Softmax}(QK^{T}/\sqrt{d})$, where input-dependent $Q$ and $K$ matrices are prone to optimizing for increased spectral norms—causing $A(X)$ to collapse into an identity matrix ($A(X)\approx I$) and creating an identity shortcut for anomalies—the FFU mitigates this issue through its input-independent architecture. The spectral filter (see $W_{freq}$ in Fig. 3) consists of learnable parameters that remain fixed during inference. During training, it learns a static spectral template representing the global energy distribution of normal patterns and cannot dynamically ”re-focus” its weights on unseen anomalous signals. Third, while self-attention incurs a quadratic computational cost of $O(N^{2}\cdot d)$ that bottlenecks memory and throughput, the FFU leverages the Fast Fourier Transform (FFT) to achieve $O(N\log N\cdot d)$ complexity.
Eq. 6 defines the core FFU operation. The input sequence $\mathbf{X}\in\mathbb{R}^{N\times d}$ is reshaped into a 2D representation $\mathbf{X}^{R}\in\mathbb{R}^{H\times W\times d}$, projected into the frequency domain via a 2D Discrete Fourier Transform (DFT), $\mathcal{F}$, and modulated through element-wise multiplication with a learnable, input-independent spectral filter $\mathbf{W}_{freq}$. The modulated features are then restored to the spatial domain using an Inverse DFT ($\mathcal{F}^{-1}$). Furthermore, we introduce a Global and Local Masking (GLM) mechanism on $W_{freq}$ to suppress shortcut learning. local-masking (LM) mechanism prevents tokens from attending to their own positions. Considering the global modeling strength of FFU and ViT, all tokens in the same sequence may share similar information, we propose to mask tokens globally and randomly. Specifically, by the convolution theorem, frequency-domain multiplication corresponds to spatial circular convolution:
$\mathbf{Y}=\text{FFU}(\mathbf{X})=\mathcal{F}^{-1}(\mathcal{F}(\mathbf{X}^{R})\odot\mathbf{W}_{freq})=\mathbf{X}^{R}\circledast k$ (6)
[…]
$\mathbf{W}^{M}_{freq}=\mathcal{F}(k\odot M_{local}\odot M_{global})$ (7)
The final formulated FFU is expressed as:
$\text{FFU}(\mathbf{X})=\mathcal{F}^{-1}(\mathcal{F}(\mathbf{X}^{R})\odot\mathbf{W}^{M}_{freq})$ (8)
@@
Image-Level Pixel-Level Image-Level Pixel-Level
@@ 3.1 Experimental Setting
Datasets We evaluated our method on four well-established datasets, including three for industrial scenarios: MVTec-AD (Bergmann et al., 2019), ViSA (Zou et al., 2022), Real-IAD (Wang et al., 2024), and one for medical scenario: Universal Medical (He et al., 2024b, a; Zhang et al., 2023, 2024). MVTec-AD comprises 15 object categories, with 3,629 normal images in the training set and 498 normal images alongside 1,982 anomalous samples in the test set. ViSA contains 12 categories, providing 8,659 normal training images and a test set with 962 normal images and 1,200 anomalous cases. Real-IAD, the largest industrial benchmark, includes 30 diverse objects, utilizing 36,645 normal images for training and 63,256 normal images combined with 51,329 anomalous instances for testing. Universal Medical consists of 13,339 normal cases for training, with a test set containing 2,514 normal cases and 4,499 abnormal cases. This dataset spans three medical imaging modalities: Brain MRI, Liver CT, and Retinal CT scans. All datasets include both image-level and pixel-level labels. Metrics For evaluation metrics, we followed protocols from (Zhang et al., 2023; He et al., 2024a; Guo et al., 2025), adopting four metrics: Area Under the Receiver Operating Curve (AUC), F1-max score, average precision (AP) for both image-level detection and pixel-level localization tasks, and Area Under the Per-Region-Overlap (AUPRO) to further evaluate the localization task. Implementation details In our experiments, the encoder is initialized with DINO pre-trained weights (Darcet et al., 2023) and kept frozen during training. Images are resized to 448×448 and then center-cropped to 392×392. We use a stable variant of the AdamW optimizer (Wortsman et al., 2023) incorporating the AMSGrad algorithm (Reddi et al., 2019), with a batch size of 32. Training iterations are set to 20,000 for all four datasets. The learning rate is initialized to 2e-3 and gradually reduced to 2e-4 via a Cosine Annealing schedule with a warm-start scheme (Loshchilov & Hutter, 2016). The model is optimized using the global hard-mining loss (Guo et al., 2025). In LRNB, we set the noise rate to 0.4 and the $r$ to 4. Similar to Dinomaly, set the number of LRNB layers to 1 for MVTec-AD and ViSA datasets, and 2 for more complex Universal Medical and Real-IAD datasets.
@@ 3.2 Evaluations
Comparison with SOTA MUAD methods We evaluate our method against seven state-of-the-art (SOTA) MUAD methods—UniAD (You et al., 2022), DiAD (He et al., 2024b), ViTAD (Zhang et al., 2023), Recontrast (Guo et al., 2024), and Dinomaly (Guo et al., 2025)—across four benchmark datasets: MVTec-AD (Bergmann et al., 2019), ViSA (Zou et al., 2022), Universal Medical (He et al., 2024b), and Real-IAD (Wang et al., 2024). Performance is quantified using both image-level metrics (I-AUC, I-AP, I-F1) and pixel-level metrics (P-AUC, P-AP, P-F1, P-AUPRO), where higher values indicate better detection capability. Experimental results are presented in Table 1, where our method outperforms comparative methods across all datasets in most metrics (The qualitative results obtained by our method can be seen in Fig. 4). On the widely adopted MVTec-AD, our method achieves SOTA overall performance, with the highest image-level metrics of 99.8/99.9/99.4, and pixel-level metrics of 98.5/71.5/70.5/95.6. On ViSA, our method consistently achieves the superior overall performance, i.e., image-level metrics of 98.7/98.9/95.9 and pixel-level metrics of 99.0/53.9/55.0/94.5. These results demonstrate that image-level performance on these two datasets is nearly 100%, making performance improvements subtle. On Universal Medical, our method attains the highest image-level metrics of 87.6/87.0/82.5, surpassing prior SOTAs by a large margin of 2.7/2.9/1.5, and achieves best or second-best pixel-level metrics of 97.1/56.0/55.6/87.3. On Real-IAD, our method produces a new SOTA result, with image-level and pixel-level performance of 90.8/88.2/81.5 and 98.9/46.6/49.7/95.0, outperforming previous SOTAs by 1.5/1.4/1.3 and 0.1/3.4/2.6/1.1. These results demonstrate strong generalization to complex medical and diverse real-world industrial scenarios.
Figure 4: Qualitative results of our Shortcutbreaker on the MVTec-AD, VisA, Universal Medical, and Real-IAD datasets. The first row displays the input images with their ground truth, while the second row presents the corresponding anomaly maps.
@@
Table 2: Ablations studies of component contributions on Real-IAD and MvTec dataset, including LRNB: low-rank noisy bottleneck, FFB: frequency filtering block, GF: global filtering and GLM: global-local masking. (%)
LRNB FFB Real-IAD MvTec-AD
GF GLM I-AUC/AP/F1 P-/AUC/AP/F1/AUPRO I-AUC/AP/F1 P-/AUC/AP/F1/AUPRO
basemodel 84.7/81.3/75.8 97.6/32.7/39.8/91.4 98.6/99.1/97.8 97.3/61.7/62.9/93.0
$\checkmark$ 88.7/86.6/79.5 98.5/40.7/46.4/93.5 99.7/99.8/99.4 98.4/70.1/69.8/94.9
$\checkmark$ 84.8/81.4/75.8 97.7/32.8/39.8/91.4 98.8/99.3/97.8 97.2/61.3/63.1/92.8
$\checkmark$ $\checkmark$ 85.3/81.9/76.0 97.7/33.5/40.4/91.5 99.1/99.5/98.3 97.3/62.1/63.4/93.0
$\checkmark$ $\checkmark$ 89.6/86.5/80.4 98.8/46.3/48.9/94.4 99.6/99.8/99.1 98.2/69.7/69.1/94.8
$\checkmark$ $\checkmark$ $\checkmark$ 90.8/88.2/81.5 98.9/46.6/49.7/95.0 99.8/99.9/99.4 98.5/71.5/70.5/95.6
Table 3: Effect of dual injection in LRNB on Real-IAD, G1 and G2 denotes the injection of LRNB before group 1 and group 2 decoder. (%)
LRNB I-AUC/AP/F1 P-/AUC/AP/F1/AUPRO
None 85.3/81.9/76.9 97.7/33.4/30.4/91.5
G1 90.2/87.5/81.1 98.9/45.2/48.5/94.8
G2 88.7/86.0/79.3 98.0/45.6/49/1/94.0
G1+G2 90.8/88.2/81.5 98.9/46.6/49.7/95.0
Table 4: Effect of global and local masking in FFB on Real-IAD, GM: global masking, LM: local masking and GLM: global-local masking. (%)
FFB I-AUC/AP/F1 P-/AUC/AP/F1/AUPRO
None 89.6/86.5/80.4 98.8/46.3/48.9/94.4
GM 89.8/86.9/80.5 98.8/46.3/49.2/94.7
LM 90.0/86.8/80.6 99.0/46.4/49.2/94.7
GLM 90.8/88.2/81.5 98.9/46.6/49.7/95.0
@@ 3.3 Ablation Studies
Overall Ablation To explore the contributions of each component in our method, including the low-rank noisy bottleneck (LRNB), global filtering (GF), and global-local masking (GLM) operations, we conduct ablation experiments as shown in Table LABEL:table2_new. The baseline is constructed following Dinomaly (Guo et al., 2025) and ViTAD (Zhang et al., 2023), which construct a baseline model with a DINO-pretrained ViT encoder and a learnable softmax attention-based ViT decoder. The effectiveness of the baseline has been proven in industrial scenarios. The results of the ablation experiments demonstrate that the proposed LRNB, GF and GLM modules all contribute to performance improvement, with their combinations further enhancing performance. Compared to the baseline model, using any single module alone achieves better performance in terms of the key indicators I-AUC and P-AUC. Among them, FFB (GF with GLM) brings moderate improvements, while LRNB shows the most significant enhancement—it notably boosts image-level and pixel-level performances with respective large margins of 5.1/6.9/5.7 and 2.3/13.9/9.9/3.6 on Real-IAD dataset and 1.1/0.7/1.6, and 1.1/8.4/6.9/1.9 on MvTec-AD dataset. These results demonstrate that the LRNB module serves as a core foundation in our method. When replacing self-attention with sole GF, the overall improvements on two datasets are subtle, however the computational complexity is reduced from $O(N^{2}\cdot d)$ to $O(N\log N\cdot d)$. Further combination with GLM yields even better results. The final integration of all three modules (LRNB + GF + GLM) delivers the best overall performance, achieving the highest image-level and pixel-level metrics on both datasets—confirming that their collaborative effect effectively enhances the model’s overall performance.
Table 3 illustrates the impact of dual injection within the LRNB framework. The results demonstrate that injecting LRNB before either the Group 1 (G1) or Group 2 (G2) decoder yields significant performance improvements. Notably, G1 injection serves as the primary contributor, achieving superior accuracy compared to G2. We attribute this to the fact that G1 injection influences all subsequent decoder blocks, whereas G2 only affects the final four blocks. Furthermore, the combination of G1 and G2 surpasses single-injection strategies at both the image and pixel levels by further enlarging the score gap between normal and abnormal cases. The corresponding training loss and anomaly score plots in Fig. 3 corroborate these findings. Table 4 evaluates Global-Local Masking (GLM). While GM or LM alone yields marginal gains, combining them significantly boosts performance. This confirms our observation that ViT features suffer from information leakage via both self-identity and long-range token correlations, necessitating a dual-masking approach. The GLM prevents the shortcuts globally and locally, thus enhancing the anomaly detection ability.
@@
Figure 5: The corresponding plots of training loss (left part) and the averaged anomaly scores (right part) of the methods in Table 3.
Table 5: Performance comparison of different bottlenecks (BNs) on Real-IAD. G1 denotes the bottleneck is only injected before group 1 decoder (%)
Bollteneck I-AUC/AP/F1 P-/AUC/AP/F1/AUPRO
None 84.7/81.3/75.8 97.6/32.7/39.8/91.4
FJ (G1) 87.6/84.5/78.3 98.3/40.1/44.8/92.3
NBD (G1) 88.5/86.0/79.2 98.5/43.3/47.1/93.5
LRNB (G1) 90.2/87.5/81.1 98.9/45.2/48.5/94.8
LRNB (G1+G2) 90.8/88.2/81.5 98.9/46.6/49.7/95.0
Table 6: Performance comparison of different modules in decoder on Real-IAD. (%)
Decoder Blocks I-AUC/AP/F1 P-/AUC/AP/F1/AUPRO
ViT 88.7/86.6/79.5 98.5/40.7/46.4/93.5
CNN+ViT 89.0/86.1/79.5 98.8/45.5/48.7/94.4
NMA+ViT 90.4/88.0/81.2 98.9/45.4/48.9/94.6
LSA 89.3/87.5/79.9 98.6/43.4/47.9/93.9
FFB (ours) 90.8/88.2/81.5 98.9/46.6/49.7/95.0
Comparison between LRNB and previous noisy bottlenecks Previous methods also proposed some noisy operations in the bottleneck to avoid identity shortcuts, such as feature jittering (FJ) in UniAD (You et al., 2022) and noisy dropout bottleneck (NDB) in Dinomaly (Guo et al., 2025). To further validate the advantage of the proposed LRNB, we conduct a comparison in Table 6. The results in this table demonstrate that FJ experiences a performance drop and NDB yields slight improvements compared to models without such noise operations, while our LRNB and LRNB-D achieve a significant enhancement.
Table 7: Comparison of computational cost of different methods on Real-IAD, mAD is the average value of all the metrics. Shortcutbreaker-s is based on the encoder of ViT-small model, Shortcutbreaker-s* is based on ViT-small model and trained with 224x224 images.
Method Params(M) FLOPS(G) mAD (%)
UniAD 24.5 3.6 67.5
DiAD 1331.3 451.5 52.6
MambaAD 25.7 8.3 72.7
Dinomaly 132.8 104.7 77.0
Shortcutbreaker 109.8 86.5 78.7
Shortcutbreaker-s 27.6 21.8 77.5
Shortcutbreaker-s* 27.6 7.19 73.3
@@
To explore the capacity of these operations in addressing the ”identity mapping” issue, we further examine the corresponding training loss and average anomaly score. As shown in Fig. 6, models without a bottleneck achieve near-zero training loss and obtain extremely close anomaly scores between normal and abnormal cases (0.024 vs. 0.029). These results indicate that these models suffer from identity mapping behavior where inputs are directly replicated in outputs. Introducing FJ or NDB noise operations partially alleviates this issue, which we attribute to enhanced noise robustness from exposure to diverse industrial patterns. Our LRNB and LRNB-D effectively resolve identity mapping: while moderately increasing the score of normal data, it substantially enhances the score of abnormal cases. This strategic trade-off widens the discrimination margin between normal and abnormal samples, ultimately improving anomaly detection performance.
Figure 6: The plots of training loss (left part) and the averaged anomaly scores (right part) of different bottlenecks on Real-IAD.
Comparison between FFB and previous modules in decoder Prior works have introduced specialized decoder modules to mitigate identity mapping. For instance, (Lu et al., 2024) attributes this issue to encoder-decoder homogeneity and advocates for heterogeneous structures, such as using CNN blocks to reconstruct ViT encoder outputs. Similarly, (You et al., 2022) introduced neighbor-masking attention (NMA) to prevent information leakage, while (Guo et al., 2025) utilized linear self-attention (LSA) to ”un-focus” features for robust reconstruction. As shown in Table 6, replacing the ViT baseline with a hybrid ViT+CNN structure improves performance, with LSA providing further gains. Additionally, NMA proves effective in the ViT decoder, yielding improvements of 0.7/1.4/1.7 in image-level and 0.1/4.7/2.5/1.0 in pixel-level metrics over pure ViT with negligible computational overhead. Ultimately, our FFB outperforms all existing methods across all metrics, which we attribute to its superior adaptability to ViT-extracted tokens.
Comparison of computational complexity Table 7 presents the trade-off between computational complexity and mAD on Real-IAD dataset. where our Shortcutbreaker framework demonstrates clear superiority over state-of-the-art methods. Notably, the primary Shortcutbreaker variant achieves a peak performance of 78.7 mAD, surpassing the second-best Dinomaly while reducing computational costs by 17.4%. Furthermore, Shortcutbreaker-S and Shortcutbreaker-S*offer exceptional scalability; the former maintains a superior 77.5 mAD with only 27.6M parameters, while the latter delivers a highly efficient 7.19G FLOPs for real-time applications, even better than the light-weight structure of MambaAD (He et al., 2024a).